OpenStudy (anonymous):

f(x)=7x+6/x+5 what is the inverse

OpenStudy (anonymous):

When finding the inverse, you exchange the x's and y's. like this: x=7y+6/y+5

OpenStudy (anonymous):

Then you solve for y.

OpenStudy (anonymous):

Is this how I do it? x=(7y+6)/(y+5)

OpenStudy (anonymous):

Yea, but when i was learning this, i was told to solve for y, or f(x) as well.

OpenStudy (anonymous):

ok my answer I got is x=6/y,2

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

but then I did f(x)=(7y+6)/(y+5) and got -5x-6/x-7

OpenStudy (anonymous):

Dont know what you did there..but i said to solve for y. Meaning everything will equal to y. But that depends, how does your teacher or proffessor like it to be in when its solved?

OpenStudy (anonymous):

did I set it up right? 7y+6/y+5

OpenStudy (anonymous):

Yes, but what is it equal to?

OpenStudy (anonymous):

i don't know what I am doin wrong

OpenStudy (anonymous):

\[x=\frac{7y+6}{y+5}\]

OpenStudy (anonymous):

you interchange the x and the y's, so wherever there is an x, switch with y and wherever there is a y, switch with x

OpenStudy (anonymous):

I thought I did that...

OpenStudy (anonymous):

ok now solve for y

OpenStudy (anonymous):

i can't fiqure it out..

OpenStudy (anonymous):

You multiply each side by (y+5) to get rid of the fraction.

OpenStudy (anonymous):

im still lost

OpenStudy (anonymous):

the other way would be multiply by y+5 \[x(y+5)=7y+6\]

OpenStudy (anonymous):

\[xy+5x=7y+6\] distribution property \[xy-7y+5x=6\] subtraction property \[xy-7y=6-5x\] subtraction propety \[y=\frac{6-5x}{x-7}\]

OpenStudy (anonymous):

i just factored the last part

OpenStudy (anonymous):

pullout out a y and divided by x-7

OpenStudy (anonymous):

ignore my first answer... it dawned on me that 1/x is a reciprocal of the equation and the inverse is not the reciprocal

OpenStudy (anonymous):

NP.. Thank you for your help...